# Do LLMs Feel? Teaching Emotion Recognition with Prompts, Retrieval, and Curriculum Learning

### v1

@@ Introduction
We design a two-stage prompt strategy for data generation. In the first stage, the model generates 30 specific subtopics under a given scenario to ensure content richness and diversity. In the second stage, the model produces contextually coherent two-person dialogues based on these subtopics, with each utterance annotated with a corresponding emotion label. By guiding the model to generate emotionally targeted content—especially for underrepresented emotions—we achieve a relatively balanced distribution of emotion samples while maintaining logical consistency within the dialogues, thereby providing a more diverse and structurally balanced corpus for downstream models.
@@
All experiments are conducted on a single NVIDIA 4090D GPU. For the IEMOCAP dataset, Qwen2.5-7B-Instruct [^1] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct is used as the base model; for the MELD dataset, Qwen3-8B [^2] https://huggingface.co/Qwen/Qwen3-8B serves as the base model. The external large language model is consistently Qwen3-14B [^3] https://huggingface.co/Qwen/Qwen3-14B, which is employed to generate explicit and implicit emotion interpretations as well as speaker characteristics. Additional details on training hyperparameters and settings are provided in Appendix B. All experiments in this paper—including comparative experiments—are repeated using five different random seeds, and the average results are reported.
@@ Conclusion
This paper proposes PRC-Emo, a novel training framework for Emotion Recognition in Conversation (ERC) that integrates Prompt engineering, demonstration Retrieval, and Curriculum learning using large language models (LLMs). PRC-Emo builds the first dedicated demonstration retrieval repository for ERC and designs emotion-sensitive prompt templates based on explicit and implicit emotional cues to better understand psychological states. The curriculum learning strategy organizes training from easy to hard based on weighted emotional shifts between same-speaker and different-speaker utterances. Experimental results on the IEMOCAP and MELD benchmark datasets show that PRC-Emo achieves new state-of-the-art performance. This work highlights the potential of combining prompt-based learning with curriculum learning strategies to advance ERC tasks.

### v3

@@ Introduction
We adopt a two-stage prompt strategy for data generation. First, the model produces 30 subtopics under a given scenario to increase diversity. Then, it generates coherent two-person dialogues for these subtopics, with each utterance assigned an emotion label. This process enables targeted creation of underrepresented emotions while preserving contextual consistency, resulting in a more balanced and diverse corpus for downstream ERC models.
@@
All experiments are conducted on a single NVIDIA 4090D GPU. For the IEMOCAP dataset, Qwen2.5-7B-Instruct is used as the base model; for the MELD dataset, Qwen3-8B serves as the base model. The external large language model is consistently Qwen3-14B, which is employed to generate explicit and implicit emotion interpretations as well as speaker characteristics. Additional details on training hyperparameters and settings are provided in Appendix B. All experiments in this paper—including comparative experiments—are repeated using five different random seeds, and the average results are reported.
@@ Conclusion
This paper proposes PRC-Emo, a novel training framework for Emotion Recognition in Conversation (ERC) that integrates Prompt engineering, demonstration Retrieval, and Curriculum learning using large language models (LLMs). PRC-Emo builds the first dedicated demonstration retrieval repository for ERC and designs emotion-sensitive prompt templates based on explicit and implicit emotional cues to better understand psychological states. The curriculum learning strategy organizes training from easy to hard based on weighted emotional shifts between same-speaker and different-speaker utterances. Experimental results on the IEMOCAP and MELD benchmark datasets show that PRC-Emo achieves new state-of-the-art performance. This work highlights the potential of combining prompt-based learning with curriculum learning strategies to advance ERC tasks. In future work, we plan to further explore stronger LLMs, more efficient prompting paradigms, and advanced curriculum designs to enhance emotional reasoning and improve the robustness and generalization of ERC systems.
Acknowledgments
This work is funded in part by the National Natural Science Foundation of China Project (No. 62372078).